# Sub-GeV Dark Matter Direct Detection with Neutrino Observatories

### v2

@@ Abstract
We present a new technique for sub-GeV dark matter (DM) searches and a new use of neutrino observatories. DM-electron scattering in an observatory can excite or ionize target molecules, which then produce light that can be detected by the photomultiplier tubes (PMTs). While individual DM scatterings are undetectable, the aggregate rate from many independent scatterings can be isolated from the total PMT dark rate using the expected DM annual modulation. We showcase this technique with the example of JUNO, a 20,000-ton scintillator detector, showing that its sensitivity in some mass ranges exceeds all other techniques and reaches key particle-theory benchmarks.
@@
In this Letter, we present a new technique for sub-GeV DM searches and a new use of neutrino observatories. The basic ideas are as follows. DM-electron scattering in an observatory excites or ionizes target molecules, producing eV-range photons that can be detected with the photomultiplier tubes (PMTs). Individual DM-electron scattering events are undetectable because they typically lead to less than one PMT hit. However, we show that neutrino observatories can search for the aggregate signal rate from many independent DM scattering events. This signal rate must be compared to the observatory’s dark rate, which is the total rate of PMT hits due to thermal noise and other sources. Though the DM signal rate is small compared to the dark rate, it can be isolated using the annual modulation due to Earth’s orbital motion. This is a new use of neutrino observatories, which use many PMT hits per event to reconstruct signals and backgrounds. Figure 1 illustrates the new DM signal.
We demonstrate this new technique through the example of the Jiangmen Underground Neutrino Observatory (JUNO) [42, 43]. The first advantage is a huge target volume — here 20,000 ton — whereas direct-detection experiments have only reached the ton scale. The second is that the detector material — linear alkylbenzene (LAB), a liquid scintillator — has an electronic excitation threshold of only a few eV, whereas noble liquids have thresholds of order 10–100 eV. The disadvantage of neutrino observatories compared to direct-detection experiments is much larger backgrounds. However, we show that the net effect is that JUNO has sensitivity to DM-electron interactions in some mass ranges that (A) exceeds all other techniques and (B) reaches some important theoretical targets. Although we focus on JUNO, our technique can be adapted to other neutrino observatories. This new use of neutrino observatories complements familiar techniques, $e.g.$, searches for DM annihilation products [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] and the scattering of boosted [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77], heavy, or strongly interacting DM [78, 79, 80, 81, 82, 83, 84].
Figure 1: Schematic of the new technique. Multiple independent DM scatterings with molecules in a neutrino observatory produce photons that can be detected by the PMTs.
@@
JUNO’s dark rate, dominated by thermal noise from the PMTs themselves, will be high [86, 87]. Its 20-inch PMTs trigger at about 25 kHz each, while its 3-inch PMTs trigger at about 5 kHz each, leading to a total dark rate rate of about $6\times 10^{5}$ kHz, $i.e.$, about $5\times 10^{13}$ PMT hits per day or $2\times 10^{16}$ per year. Those rates were obtained at surface temperatures and background levels, plus after only short periods in darkness. Once the PMTs are deployed in JUNO, all of those factors will be more favorable to lower rates. For our purposes, it will be enough for JUNO to track the total number of hits per small time interval, which we take to be a day.
@@
The high speeds of bound electrons relative to DM enhance the likelihood of excitations sufficient to overcome this threshold. For LAB (with an excitation threshold of about 5 eV), this means that excitation largely proceeds for DM with masses above about 10 MeV.
@@
$(\sigma v)_{ij}=\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}\,(2\pi)\delta(\Delta E_{ij}-\omega)\frac{\absolutevalue{\mathcal{M}_{\text{free}}}^{2}\absolutevalue{f_{ij}(\vec{q})}^{2}}{16m_{\chi}^{2}m_{e}^{2}},$ (3)
@@
|$\frac{\absolutevalue{\mathcal{M}_{\text{free}}}^{2}}{16m_{\chi}^{2}m_{e}^{2}}\equiv\frac{\pi\bar{\sigma}_{e}}{\mu_{\chi e}^{2}}F_{\text{DM}}^{2}(q),$ (4)
@@
To the best of our knowledge, LAB has not been considered as a sub-GeV DM detection material, but similar aromatic organic targets have been, $e.g.$, EJ-301 [22], which also has a benzene ring. In Ref. [22], the form factor of EJ-301 was simply taken to be that for benzene, as the $\pi$-electrons in the benzene ring are expected to dominate the rate. While the lowest-energy transition is suppressed for benzene alone (its symmetry imposes selection rules that suppress some electronic transitions), it will not be for LAB, because the long alkyl chain attached to the benzene ring lowers the symmetry of the molecule, changing the form factor. To be conservative, we also use the benzene form factor alone [22]. We expect that a proper treatment of the LAB form factor will improve sensitivity. We will do this in future work.
@@
$g_{\chi}(v)=\frac{1}{K}\exp\left(-\frac{\absolutevalue{\vec{v}+\vec{v_{E}}}^{2}}{v_{0}^{2}}\right)\Theta(v_{\text{esc}}-\absolutevalue{\vec{v}+\vec{v_{E}}})\,,$ (6)
@@
$R_{\rm sig}=\xi\frac{N^{e}_{\rm LAB}\rho_{\chi}\bar{\sigma}_{e}}{8\pi m_{\chi}\mu_{\chi e}^{2}}\sum_{i,j}\int\frac{d^{3}\vec{q}}{q}\,\eta(v^{ij}_{\text{min}}(q))F_{\text{DM}}^{2}(q)\absolutevalue{f_{ij}(\vec{q})}^{2},$ (9)
@@
DM Sensitivity from Annual Modulation.— In many dedicated DM detectors, the backgrounds are small after cuts, $e.g.$, requiring that events arise within a fiducial volume, occur in a target energy range, or have certain characteristics. Here, such cuts are not possible because the only observable is the total rate of PMT hits. This rate is dominated by the background (the dark rate, $B\sim 2\times 10^{16}$ events/yr [86, 87]). Because the background cannot be predicted precisely, we cannot estimate the DM signal sensitivity by requiring $S_{\rm tot}\gtrsim\sqrt{B_{\rm tot}}$, where those are the total numbers of signal and background events in the observation period, $T$.
However, we can isolate even a small DM signal through the annual modulation caused by Earth’s orbital motion [93, 94]. The modulation amplitude, $f_{\text{mod}}$, is determined via the change in the scattering rate that appears via Eq. (9). The rate as written is used on 2 March and 2 September, while it is modified to take into account the fact that the DM average velocity increases by 15 km/s on 2 June and decreases by 15 km/s on 2 December. We find that $f_{\text{mod}}\sim 0.1$; it is somewhat larger (smaller) for light (heavy) DM, which we take into account. The DM signal is then
@@
Figure 2: JUNO’s annual-modulation $95\%$ C.L. sensitivity to the DM-electron scattering cross section as a function of DM mass, assuming a nominal one year of exposure (20,000 ton-yr). We also show estimated one-year sensitivities for Borexino and SNO+. We show results for a heavy mediator (left panel) and a light mediator (right panel), along with the complementary direct detection [95, 96, 97, 98, 99, 100] and solar reflection (“Solar”) [101, 102] constraints. “Theory Benchmark” shows parameter spaces for some benchmark particle models [1]. Note differing axes between the two panels.
@@
To isolate a DM signal, one should fit the binned time-series data for $N(t)=B(t)+S(t)$. For the case where the background rate is constant, $B(t)=B_{\text{tot}}/T$, the basic results and uncertainties can be obtained via simple methods [103, 104, 105]. The modulated part of the signal counts is $S_{\text{tot}}f_{\text{mod}}$ and the uncertainty on the background counts is $\sqrt{B_{\text{tot}}}$. To set the 95% C.L. sensitivity, we therefore use $S_{\text{tot}}=2\,\sqrt{B_{\text{tot}}}/f_{\text{mod}}\sim 20\sqrt{B_{\text{tot}}}$, where we take into account how $f_{\text{mod}}$ varies with mass. There is no trials factor in searching for this DM signal modulation, as its amplitude, period, and phase are all known.
@@
In real experiments, the raw measured dark rates are not constant (see, $e.g.$, Borexino [106], Super-K [107], and IceCube [108]). There are secular trends as well as short-timescale fluctuations. Careful work by JUNO will be needed to pre-process the data to avoid reaching spurious conclusions [109]. One step should be removing all physics events above $\sim$0.5 MeV and flashers, both of which produce correlated hits across multiple PMTs, as well as anomalous activity of individual PMTs. While this cleaning will not appreciably reduce the dark rate, it will greatly reduce its non-thermal-noise components. A second step should be using independent measurements of detector properties to detrend obscuring variations in the data. In JUNO, a powerful tool for this will be the dark rate of the PMTs in the cylindrical, optically isolated outer detector [110], which is filled with water, for which DM signals are expected to be much smaller. Ultimately, we expect that several years of data will be needed for robust conclusions, though more conservative analyses could be done sooner, including by using subsets of PMTs.
@@
Once JUNO obtains a cleaned, detrended total rate, non-statistical fluctuations may remain an obstacle. Those that occur on timescales much less than one year will average out. Of greater concern are components of the background that have annual modulations, some of which are discussed in Ref. [111]. PMT thermal noise depends on the detector temperature, radon and muons depend on atmospheric conditions, solar neutrinos depend on Earth’s distance to the Sun, and reactor antineutrinos depend on power cycles. The cleaning and detrending procedures discussed above should greatly reduce the effects of all of these. In addition, independent measurements will be possible. In any case, these modulations are shifted in phase from the DM signal by weeks to months and will have small phase uncertainties. Finally, combining the results of multiple detectors (discussed below) would make the analysis more robust, as the signals depend on the detector properties, while the backgrounds depend on the detector properties and location.
@@
Figure 2 shows the projected sensitivity to the DM-electron scattering cross section as a function of DM mass. We show results for two benchmark scenarios for the DM-electron interaction: via a heavy mediator ($F_{\text{DM}}=1$) and via a light mediator ($F_{\text{DM}}(q)=(\alpha m_{e})^{2}/q^{2}$). We also show complementary constraints from direct-detection experiments and solar reflection. The orange shaded regions show parameter spaces for some concrete benchmark models that produce the observed relic DM abundance, though the parameter space of viable DM models is of course broader.
@@
Figure 2 also shows simple estimates of the one-year sensitivities with some example liquid scintillator neutrino experiments (Borexino and SNO$+$). As an approximation, we assumed the same PMT efficiencies, in-medium spectrum, and propagation as for JUNO, though these inputs will likely be somewhat worse. We also assumed the same form factor as for JUNO, as these are all benzene-based scintillators. We updated the calculation for their specific volumes, dark rates, and PMT coverage. This shows that this technique is promising for multiple detectors.
@@
The position of the minimum near 10 MeV in Fig. 2 is driven by the low-excitation threshold of LAB being a few eV, as discussed earlier. At higher masses, the sensitivity degrades linearly due to the lower DM number density. At lower masses, it degrades rapidly because fewer DM particles have sufficient momentum to overcome the excitation threshold. Because the modulation fraction grows with decreasing DM mass, the minimum is shifted to slightly lower DM masses (in agreement with, $e.g.$, Ref. [5]).
@@
We introduce a new technique to probe sub-GeV DM and a new use of neutrino observatories. Specifically, we demonstrate that DM-electron scattering in neutrino observatories can produce detectable signals through the excitation of target molecules. While individual scattering events are undetectable, the aggregate rate from many independent scatters can be isolated from the PMT dark rate through annual modulation. Focusing on JUNO, we show that this method enables world-leading sensitivity to the DM-electron scattering cross section in some mass ranges, reaching cross sections as low as a few times $10^{-39}$ cm2 for MeV-scale DM with a nominal one year of data. This power comes from JUNO’s huge size and high fluorescence quantum yield.
Next steps include refinement of the theoretical and experimental details for this technique in JUNO. This requires a deeper understanding of the detector’s background rates, improved modeling of the signal response, and a detailed study of potential systematic uncertainties. With JUNO’s data taking starting in 2025, early-stage experimental validation of key inputs would accelerate these steps. While there will be experimental challenges in implementing our new technique, we are optimistic that they can be surmounted, given the importance of the task.
Beyond JUNO, this new technique can be generalized to other neutrino observatories, both scintillator-based (those noted above plus Daya Bay, KamLAND, and more) and otherwise (Super-K/Hyper-K, IceCube, and DUNE). It would also be interesting to explore the potential of this technique for dedicated DM detectors. Utilizing multiple detectors is important because they will have very different backgrounds and systematics. Further, if a signal were observed in JUNO, they would extract only a degenerate combination of DM mass and cross section — a characteristic trajectory in Fig. 2, whereas other detectors would have different trajectories, which could cross JUNO’s. In future work, we will address these and other topics.

### v3

@@ Abstract
We present a new technique for sub-GeV dark matter (DM) searches and a new use of neutrino observatories. DM-electron scattering in an observatory can excite or ionize target molecules, which then produce light that can be detected by the photomultiplier tubes (PMTs). While individual DM scatterings are indistinguishable, the aggregate rate from many independent scatterings can be isolated from the total PMT dark rate using the expected DM annual modulation. We showcase this technique with the example of JUNO, a 20,000-ton scintillator detector, showing that its potential sensitivity in some mass ranges exceeds other techniques and reaches key particle-theory benchmarks.
@@
In this Letter, we present a new technique for sub-GeV DM searches and a new use of neutrino observatories. The basic ideas are as follows. DM-electron scattering in an observatory excites or ionizes target molecules, producing eV-range photons that can be detected with the photomultiplier tubes (PMTs). Individual DM-electron scattering events are indistinguishable because they typically lead to less than one PMT hit. However, we show that neutrino observatories can search for the aggregate single-hit signal rate from many independent DM scattering events. This signal rate must be compared to the observatory’s dark rate, which is the total rate of PMT hits due to thermal noise and other sources. Though the DM signal rate is small compared to the dark rate, it can be isolated using the annual modulation due to Earth’s orbital motion. This is a new use of neutrino observatories, which use many PMT hits per event to reconstruct signals and backgrounds. Figure 1 illustrates the new DM signal.
We demonstrate this new technique through the example of the Jiangmen Underground Neutrino Observatory (JUNO) [42, 43]. The first advantage is a huge target volume — here 20,000 ton — whereas direct-detection experiments have only reached the ton scale. The second is that the detector material — linear alkylbenzene (LAB), a liquid scintillator — has an electronic excitation threshold of only a few eV, whereas noble liquids have thresholds of order 10–100 eV, due to differences in their electronic structure. The disadvantage of neutrino observatories compared to direct-detection experiments is much larger backgrounds. However, we show that the net effect is that JUNO has sensitivity to DM-electron interactions in some mass ranges that (A) exceeds all other techniques and (B) reaches some important theoretical targets. Although we focus on JUNO, our technique can be adapted to other neutrino observatories; combining several results will make it even more robust. This new use of neutrino observatories complements familiar techniques, $e.g.$, searches for DM annihilation products [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61] and the scattering of boosted [62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], heavy, or strongly interacting DM [79, 80, 81, 82, 83, 84, 85].
Figure 1: Schematic of the new proposal. The aggregate rate of uncorrelated DM scatterings with molecules in a neutrino observatory produces a statistical signal due to photons detected by the PMTs.
@@
JUNO’s dark rate, dominated by thermal noise from the PMTs themselves, will be high [87, 88]. Its 20-inch PMTs trigger at about 25 kHz each, while its 3-inch PMTs trigger at about 5 kHz each, leading to a total dark rate rate of about $6\times 10^{5}$ kHz, $i.e.$, about $5\times 10^{13}$ PMT hits per day or $2\times 10^{16}$ per year. For our purposes, it will be enough for JUNO to track the total number of hits per small time interval, which we take to be a day.
@@
The high speeds of bound electrons relative to DM enhance the likelihood of excitations sufficient to overcome this threshold. For LAB (with an excitation threshold of about 5 eV), this corresponds to a typical minimum mass of about 10 MeV, above which most DM particles in the halo have sufficient kinetic energy to excite electrons. Below roughly this mass, only particles in the higher-velocity tail of the distribution contribute, making the excitation rate increasingly suppressed.
@@
[…]
@@
[…]
@@
To the best of our knowledge, the low-excitation threshold properties of LAB have not been exploited for sub-GeV DM detection, but similar aromatic organic targets have, $e.g.$, EJ-301 [22], which also has a benzene ring. In Ref. [22], the form factor of EJ-301 was simply taken to be that for benzene, as the $\pi$-electrons in the benzene ring are expected to dominate the rate. While the lowest-energy transition is suppressed for benzene alone (its symmetry imposes selection rules that suppress some electronic transitions), it will not be for LAB, because the long alkyl chain attached to the benzene ring lowers the symmetry of the molecule, changing the form factor. This symmetry breaking lifts the suppression of the lowest-energy excitation, enhancing the response at low energies. Therefore, taking the benzene form factor underestimates the rate at low momentum transfer, and is a conservative choice. We also adopt this approach in our analysis, using the benzene form factor alone [22]. We expect that a proper treatment of the LAB form factor will improve sensitivity. We will do this in future work.
@@
where $v_{E}=240\text{ km/s}$ is taken as the typical speed of Earth, $v_{0}=230\text{ km/s}$ is the local mean speed, and $v_{\text{esc}}=600\text{ km/s}$ is the Galactic escape speed. Last, $K$ is a normalization factor, obtained from enforcing the integral of Eq. (6) over $d^{3}\vec{v}$ to be unity. We also define [5, 22]
@@
[…]
@@
DM Sensitivity from Annual Modulation.— In many dedicated DM detectors, the backgrounds are small after cuts, $e.g.$, requiring that events arise within a fiducial volume, occur in a target energy range, or have certain characteristics. Here, such cuts are not possible because the only observable is the total rate of PMT hits. This rate is dominated by the background (the dark rate, $B\sim 2\times 10^{16}$ events/yr [87, 88]). A conservative DM bound could be set by requiring that the signal rate is less than this background. Because the background cannot be predicted precisely, we cannot estimate the DM signal sensitivity by requiring $S_{\rm tot}\gtrsim\sqrt{B_{\rm tot}}$, where those are the total numbers of signal and background events in the observation period, $T$.
However, we can isolate even a small DM signal through the annual modulation caused by Earth’s orbital motion [94, 95]. The modulation amplitude, $f_{\text{mod}}$, is determined via the change in the scattering rate that appears via Eq. (9). The rate as written is used on 2 March and 2 September, while it is modified to take into account the fact that the Earth velocity increases by 15 km/s on 2 June and decreases by 15 km/s on 2 December. We find that $f_{\text{mod}}\sim 0.1$; it is somewhat larger (smaller) for light (heavy) DM, which we take into account. The DM signal is then
@@
Figure 2: JUNO’s annual-modulation potential $95\%$ C.L. sensitivity for the DM-electron scattering cross section as a function of DM mass, assuming a nominal one year of exposure (20,000 ton-yr). We also show estimated potential one-year sensitivities for Borexino and SNO+. We show results for a heavy mediator (left panel) and a light mediator (right panel), along with the complementary direct detection [96, 97, 98, 99, 100, 101] and solar reflection (“Solar”) [102, 103] constraints. “Theory Benchmark” shows parameter spaces for some benchmark particle models [1]. Note differing axes between the two panels.
@@
To isolate a DM signal, one should fit the binned time-series data for $N(t)=B(t)+S(t)$. For the case where the background rate is constant, $B(t)=B_{\text{tot}}/T$, the basic results and uncertainties can be obtained via simple methods [104, 105, 106]; we discuss in the next two paragraphs and Supplemental Material the challenges in real data. The modulated part of the signal counts is $S_{\text{tot}}f_{\text{mod}}$ and the uncertainty on the background counts is $\sqrt{B_{\text{tot}}}$. To set the 95% C.L. sensitivity, we therefore use $S_{\text{tot}}=2\,\sqrt{B_{\text{tot}}}/f_{\text{mod}}\sim 20\sqrt{B_{\text{tot}}}$, where we take into account how $f_{\text{mod}}$ varies with mass. There is no trials factor in searching for this DM signal modulation, as its amplitude, period, and phase are all known. Given the background rate of $\sim 2\times 10^{16}$ events/yr, this leads to a minimum DM signal rate of $S_{\rm tot}\sim 3\times 10^{9}$ events/yr.
@@
In real experiments, the raw measured dark rates are not constant (see, $e.g.$, Borexino [107], Super-K [108], and IceCube [109]). There are secular trends as well as short-timescale fluctuations. Careful work by JUNO will be needed to pre-process the data to avoid reaching spurious conclusions [110]. One step should be removing all physics events above $\sim$0.5 MeV and flashers, both of which produce hits across multiple PMTs correlated to one physics event, as well as anomalous activity of individual PMTs. While this cleaning will not appreciably reduce the dark rate, it will greatly reduce its non-thermal-noise components, though residuals can remain [109]. A second step should be using independent measurements of detector properties to detrend obscuring variations in the data. In JUNO, a powerful tool for this will be the dark rate of the PMTs in the cylindrical, optically isolated outer detector [111], which is filled with water, for which DM signals are expected to be much smaller. Ultimately, we expect that several years of data will be needed for robust conclusions, though more conservative analyses could be done sooner, including by using subsets of PMTs. Note that we require only the scalar PMT hit rate, $i.e.$, the number of PMT hits per unit time, which can be logged at coarse time resolution ($e.g.$, daily) and is commonly recorded in neutrino detectors for monitoring purposes. This avoids the need for full per-event readout, and imposes negligible bandwidth requirements. JUNO’s existing infrastructure is expected to support this without hardware changes.
@@
Once JUNO obtains a cleaned, detrended total rate, non-statistical fluctuations may remain an obstacle. Those that occur on timescales much less than one year will average out. Of greater concern are components of the background that have annual modulations, some of which are discussed in Refs. [112, 113, 114]. PMT thermal noise depends on the detector temperature, radon and muons depend on atmospheric conditions, solar neutrinos depend on Earth’s distance to the Sun, and reactor antineutrinos depend on power cycles. The cleaning and detrending procedures discussed above should greatly reduce the effects of all of these. In addition, independent measurements will be possible. In any case, these modulations are shifted in phase from the DM signal by weeks to months and will have small phase uncertainties. Finally, combining the results of multiple detectors (discussed below) would make the analysis more robust, as the signals depend on the detector properties, while the backgrounds depend on the detector properties and location. Further details are given in Supplemental Material.
@@
Figure 2 shows the potential sensitivity (i.e., the theoretical sensitivity, which will take experimental work to approach) for the DM-electron scattering cross section as a function of DM mass. We show results for two benchmark scenarios for the DM-electron interaction: via a heavy mediator ($F_{\text{DM}}=1$) and via a light mediator ($F_{\text{DM}}(q)=(\alpha m_{e})^{2}/q^{2}$). We show results for one year of exposure; over longer periods, additional sensitivity is gained as square root of exposure time. We also show complementary constraints from direct-detection experiments and solar reflection. The relative performance of our neutrino bounds versus direct detection experiments differs between panels due to the distinct energy dependence of the light and heavy mediator interactions, and the differing thresholds of the experiments. The orange shaded regions show parameter spaces for some concrete benchmark models that produce the observed relic DM abundance, though the parameter space of viable DM models is of course broader, see, $e.g.$, Refs. [115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 3, 137, 138, 139, 140, 141, 142, 4, 143, 144, 145, 146, 147, 5, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187].
@@
Figure 2 also shows simple estimates of potential one-year sensitivities with some example liquid scintillator neutrino experiments (Borexino and SNO++). As an approximation, we assumed the same PMT efficiencies, in-medium spectrum, and propagation as for JUNO, though these inputs will likely be somewhat worse. We also assumed the same form factor as for JUNO, as these are all benzene-based scintillators. We updated the calculation for their specific volumes, dark rates, and PMT coverage. This shows that this technique is promising for multiple detectors. As the cleaned and detrended data are not publicly available, these are only projections.
@@
The position of the minimum near 10 MeV in Fig. 2 is driven by the low-excitation threshold of LAB being a few eV, as discussed earlier. At higher masses, the sensitivity degrades linearly due to the lower DM number density. At lower masses, it degrades rapidly because fewer DM particles have sufficient momentum to overcome the excitation threshold. Because the modulation fraction grows with decreasing DM mass, the minimum is shifted to slightly lower DM masses (in agreement with, $e.g.$, Ref. [5]). Note that near threshold, the modulation of the DM signal might include other detectable harmonics, which could improve signal recovery [188].
@@
We introduce a new technique to probe sub-GeV DM and a new use of neutrino observatories. Specifically, we demonstrate that DM-electron scattering in neutrino observatories can produce detectable signals through the excitation of target molecules. While individual scattering events are indistinguishable, the aggregate rate from many independent scatters can be isolated from the PMT dark rate through annual modulation. Focusing on JUNO, we show that this method enables world-leading sensitivity to the DM-electron scattering cross section in some mass ranges, reaching cross sections as low as about $10^{-39}$ cm2 for MeV-scale DM with a nominal one year of data. This power comes from JUNO’s huge size and high fluorescence quantum yield.
The goal for this first paper is to show the community that there is an exciting possible way forward but that it will be challenging, requiring substantial work from both theorists and experimentalists. Next steps include refinement of the theoretical and experimental details for this technique in JUNO. This requires a deeper understanding of the detector’s background rates, improved modeling of the signal response, and a detailed study of potential systematic uncertainties. With JUNO’s data taking starting in 2025, early-stage experimental validation of key inputs would accelerate these steps. While there will be experimental challenges in implementing our new technique, we are optimistic that they can be surmounted, given the importance of the task.
Beyond JUNO, this new technique can be generalized to other neutrino observatories, both scintillator-based (those noted above plus Daya Bay, KamLAND, and more) and otherwise (Super-K/Hyper-K, IceCube, and DUNE). It would also be interesting to explore the potential of this technique for dedicated DM detectors. Utilizing multiple detectors is important because they will have very different backgrounds and systematics; combining results would be more robust. Further, if a signal were observed in JUNO, they would extract only a degenerate combination of DM mass and cross section — a characteristic trajectory in Fig. 2, whereas other detectors would have different trajectories, which could cross JUNO’s. In future work, we will address these and other topics.